# Transformer Layers as Painters

### v3

@@ 1 Introduction
To answer these questions we perform a series of experiments on pretrained LLMs. These include experimenting with variations on the standard transformer execution strategy, and measuring the impact of these variations on the models’ performance across a variety of benchmarks for both decoder-only (Llama) and encoder-only (BERT) models. Note that our experiments never involve finetuning or otherwise adjusting the models’ parameters (with the caveat that the GLUE evaluation standard procedure includes a finetuning step for our BERT-Large model) 111 Our code is available at https://github.com/floatingbigcat/transformer-as-painter .
Figure 1: Different execution strategies.
@@ 5 Discussion
For other future work, we plan to “thaw” our models and investigate if (and how long) transformers take to adjust to the variations in the paper via fine-tuning. If these models were actually fine-tuned with their new architecture, performance would probably be even better. It’s worth noting that Parallel and Skip both have potentially lower latencies than the full model (assuming enough memory to execute the layers simultaneously). For example, the latency for the Parallel Layer for Llama2-7B for N=8 should be about half that of normal Llama2-7B. Though the aim of this paper is to better understand layers in transformer-based LLMs as opposed to introducing new models, our results suggest simple methods for easily trading accuracy for latency gains. Our results also suggest that a routing mechanism for executing frozen layers may be used for this, somewhat analogous to Switch Transformers (Fedus, Zoph, and Shazeer 2022).

### v4

@@ Abstract
Codehttps://github.com/floatingbigcat/transformer_layers_as_painters
@@ 1 Introduction
To answer these questions we perform a series of experiments on pretrained LLMs. These include experimenting with variations on the standard transformer execution strategy, and measuring the impact of these variations on the models’ performance across a variety of benchmarks for both decoder-only (Llama) and encoder-only (BERT) models. Note that our experiments never involve finetuning or otherwise adjusting the models’ parameters (with the caveat that the GLUE evaluation standard procedure includes a finetuning step for our BERT-Large model)
@@ 5 Discussion
We also plan to “thaw” our models and investigate if transformers take to adjust to the variations in the paper via fine-tuning. If these models were fine-tuned with new architectures, the performance would probably be even better. It is worth noting that Parallel and Skip both have potentially lower latencies than the full model (assuming enough memory to execute the layers simultaneously). For example, the latency for the Parallel Layer for Llama2-7B for N=8 should be about half that of normal Llama2-7B. Though the aim of this paper is to better understand layers in transformer-based LLMs as opposed to introducing new models, our results suggest simple methods to easily trade accuracy for latency gains. Our results also suggest that a routing mechanism for executing frozen layers may be used here, analogous to Switch Transformers (Fedus, Zoph, and Shazeer 2022).